# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ V Third- and higher-order interference of light
As mentioned above, the third- and higher-order interference of light is an natural extension of the second-order interference of light. The method introduced above can also be employed to calculate the third- and higher-order interference of light.
The studies of third- and higher-order interference dates back to 1963 when Glauber gave a unified definition of $N$th-order coherence function of light Glauber (1963a, c). Greenberg et al. generalized entangled photon pair to three-photon entanglement Greenberger et al. (1989, 1990). It is difficult to generate entangled three-photon directly. Bouwmeester et al. employed two entangled photon pairs to generate entangled three-photon by post-selection Bouwmeester et al. (1999). Mitchell et al. employed a laser photon and an entangled photon pair to generate entangled three-photon and observed three-photon subwavelength interference Mitchell et al. (2004). In 2010, Hübel et al. employed cascade periodically poled nonlinear crystals to generate entangled three-photon directly Hübel et al. (2010). Based on the work of Bouwmeester et al. Bouwmeester et al. (1999), Walther et al. observed four-photon subwavelength interference with two entangled photon pairs Walther et al. (2004) and Pan’s group observed as high as 12-photon subwavelength interference Zhao et al. (2004); Pan et al. (2012); Zhong et al. (2018).
Ghost imaging was originally realized with entangled photon pairs Pittman et al. (1995) and later with thermal light Gatti et al. (2004); Cheng and Han (2004); Cai and Zhu (2005); Valencia et al. (2005) via the second-order interference of light. Ghost imaging via the third- and higher-order interference of light was discussed for entangled three-photon Wen et al. (2007) and thermal light Ou and Kuang (2007); Bai and Han (2007); Agafonov et al. (2008); Zhou et al. (2010); Cao et al. (2008); Chan et al. (2009); Liu et al. (2009).
In the following, third-order interference of light is taken as an example to show how to employ Feyman’s path integral theory to calculate the third- and high-order interference of light.
@@ V.1 Third-order interference of one light beam
Figure 27 shows the third-order HBT interferometer by analogy of HBT interferometer shown in Fig. 11. The light emitted by light source (S) is split into two via a 1:2 non-polarizing beam splitter BS1. The transmitted light beam is split into two via a 1:1 non-polarizing beam splitter. The intensities detected by D1, D2, and D3 are equal. TCC is short for three-photon coincidence count and is usually written as CC for convenience. The distance between the source and detector planes are equal.
Figure 27: Third-order HBT interferometer. TCC is short for three-photon coincidence count and is usually written as CC for convenience. The meanings of other symbols are the same as the one in HBT interferometer.
Assume S is thermal light source in Fig. 27. There are six different paths to trigger a three-photon coincidence count. Figure 28 shows the six different ways. For instance, Fig. 28(a) show the path for photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3. If all the six different ways are indistinguishable, the three-photon probability function is
@@
$=$ $\langle|A_{a1,b2,c3}+A_{a1,b3,c2}+A_{a2,b1,c3}$
$+A_{a2,b3,c1}+A_{a3,b1,c2}+A_{a3,b2,c1}|^{2}\rangle.$
@@
Figure 28: Third-order HBT interferometer. TCC is short for three-photon coincidence count and is usually written as CC for convenience. The meanings of other symbols are the same as the one in HBT interferometer.
With the same method as the one for thermal light in a HBT interferometer introduced in Sect. IV.1.1, the spatial three-photon probability function is Liu and Shih (2009)
@@
where paraxial approximation and one-dimension have been employed, $\lambda$ is the wavelength of thermal light and single frequency light is assumed, $d$ is the length of one-dimension thermal light source, $L$ is the distance between the source and detector planes. Figure 29(a) shows the simulated spatial third-order coherence function of thermal light in a third-order HBT interferometer. Figure 29(b) shows the measured three-photon coincidence counts vs. the transverse distances between these three detectors. The theoretical and experimental results are consistent with each other.
Figure 29: Simulated Liu and Shih (2009) and measured Zhou et al. (2010) spatial third-order coherence functions of thermal light in the third-order interferometer.
Temporal three-photon probability function can be obtained in the same way Zhou et al. (2010)
@@
The degree of third order coherence, $g^{(3)}(0)$, of thermal light equals 6, which means that the probability to detect three photons when these three detectors are in the same space-time coordinate is 6 times of the one when these three detectors are in different coherence volumes of thermal light. It is three-photon bunching of thermal light. Three-photon superbunching can be observed with superbunching pseudothermal light Zhou et al. (2019a) and the degree of $N$th-oder coherence of thermal light equals $N!$ Liu and Shih (2009).
The third-order interference of single-mode continuous-wave laser light in the third-order HBT interferometer is simple. All the six different paths shown in Fig. 28 to trigger a three-photon coincidence count are identical. The superposition of these six probability amplitudes will not give three-photon interference. The reason is the same as the one for there is no two-photon interference for single-mode continuous-wave laser light in a HBT interferometer. No three-photon bunching can be observed when single-mode continuous-wave laser light in the third-order interferometer.
@@ V.2 Third-order interference of two independent light beams
Figure 30 shows the scheme for the third-order interference of two independent light beams. Assume S1 is a single-photon source and S2 is a single-mode continuous-wave laser. There are two different ways to trigger a three-photon coincidence count in Fig. 30. The first way is one photon is emitted by S1 and two photons are emitted by S2. The probability is $3I_{1}I_{2}^{2}/(3I_{1}I_{2}^{2}+I_{2}^{3})$, where $I_{1}$ and $I_{2}$ are the intensities emitted by S1 and S2, respectively. There are three different paths and the corresponding three-photon probability amplitudes are $A_{11,22,23}$, $A_{12,21,23}$, and $A_{13,21,22}$, respectively. The second way is these three photons are all emitted by S2. The probability is $I_{2}^{3}/(I_{1}I_{2}^{2}+I_{2}^{3})$. There is one path and the corresponding three-photon probability is $A_{21,22,23}$.
Figure 30: Scheme for the third-order interference of two independent light beams.
@@
The initial phases of photons emitted by S1 are random and emitted by S2 are identical within one coherence volume, Eq. (96) can be simplified as
@@
There is third-order interference pattern. The visibility of the third-order interference pattern is dependent on the ratio between the intensities of light emitted by single-photon source and single-mode continuous-wave laser.
@@ V.3 Third-order interference of multiple independent light beams
Figure 31 shows the scheme for the third-order interference of three independent light beams. Assume S1, S2, and S3 are three identical and independent single-photon sources. The intensities, polarizations, and the frequencies are the same for the photons emitted by these three sources. There are six different paths to trigger a three-photon coincidence count. The corresponding three probability amplitudes are $A_{11,22,33}$, $A_{11,23,31}$, $A_{12,21,33}$, $A_{12,23,31}$, $A_{13,22,31}$, and $A_{13,21,32}$, where $A_{11,22,33}$ is the probability amplitude for photon emitted by S1 is detected by D1, photon emitted by S2 is detected by D2, and photon emitted by S3 is detected by D3. The meanings of other symbols are defined similarly.
Figure 31: Scheme for the third-order interference of three independent light beams
If six different paths are indistinguishable, three-photon probability function is
@@
$=$ $\langle|A_{11,22,33}+A_{11,23,31}+A_{12,21,33}$
$+A_{12,23,31}+A_{13,22,31}+A_{13,21,32}|^{2}\rangle$
$=$ $\langle|e^{i(\varphi_{1}+\varphi_{2}+\varphi_{3})}(K_{11}K_{22}K_{33}+K_{11}K_{23}K_{31}+K_{12}K_{21}K_{33}$
$+K_{12}K_{23}K_{31}+K_{13}K_{22}K_{31}+K_{13}K_{21}K_{32})|^{2}\rangle,$
@@
With the same method and approximations above, one-dimension spatial three-photon probability function is
@@
Figure 32 shows the third-order interference pattern given by Eq. (V.3).
Figure 32: Third-order interference pattern of three independent single-photon states.
Similar method can be employed to calculate the third-order interference other three types of light beams and higher-order interference of light.
@@ VI Quantum atomic coherence theory based on Feynman’s path integral
The history of atomic coherence traces back to de Broglie in 1924 when he proposed matter wave assumption De Broglie (1924) by analogy of Planck’s Planck (1901) and Einstein’s Einstein (1905) quanta assumption of light. The de Broglie wavelength of matter wave equals
@@
De Broglie’s assumption was soon verified by Davisson et al. and Thomson et al. with diffraction pattern of electrons Davisson and Germer (1927); Thomson and Reid (1927). Due to the very short de Broglie wavelength of matter wave, it is difficult to observe the first-order interference pattern like the one of light. In 1953, Marton et al. observed first-order interference pattern of electrons in a Mach-Zehnder interferometer Marton et al. (1953). Then the first-order interference pattern of neutron Maier-Leibnitz and Springer (1962), atom Keith et al. (1991), molecule Arndt et al. (1999) were also reported. The first-order interference pattern of two independent Bose Einstein Condensations (BEC) was also reported Andrews et al. (1997).
The first second-order interference experiments with matter wave was done by Yasuda and Shimizu, in which they observed two-atom bunching with cold atoms Yasuda and Shimizu (1996). Burt et al. observed the degree of third-order coherence of cold atoms and BEC, finding that the ratio between these two equals $7.4\pm 2.6$ Burt et al. (1997). The second-order interference of fermions was also reported by several groups Henny et al. (1999); Oliver et al. (1999); Iannuzzi et al. (2006); Rom et al. (2006); Bocquillon et al. (2013).
Quantum atomic coherence theory is mainly based on Glauber’s quantum optical coherence theory Adams et al. (1994); Meystre (2001); Cronin et al. (2009). Feynman himself introduced how to employ Fenman’s path integral to calculate the first-order interference of electrons in a Young’s double-slit interferometer Feynman et al. (2010, 2011). Storey and Cohen-Tannoudji provided a good introduction about how to employ Feynman’s path integral to calculate the first-order interference of atoms Storey and Cohen-Tannoudji (1994). Employing Feynman’s path integral to calculate the first-order interference of matter wave can be found in Refs. Deng (2006, 2008); Gondran and Gondran (2005), too.
This section will summarize how to employ Feynman’s path integral to calculate the first-, second, and higher-order interference matter waves, with special attentions on the analogy between cold atoms (BEC) and thermal light (laser light).
@@ VI.1.1 First-order interference of one matter wave beam
Figure 33 shows the scheme for the first-order interference of one matter wave beam. S is a particle source. P1 and P2 are two pinholes with equal size. O is the observation plane. For simplicity, free atoms are assumed.
Figure 33: Scheme for the first-order interference of one matter wave beam.
There are two different paths to trigger a particle detection event in the observation plane. The first one is the particle going through P1. The probability is 1/2 and the probability amplitude is $A_{S1x}$. The second path is the particle going through P2. The probability is 1/2 and the probability amplitude is $A_{S2x}$. If these two different paths are indistinguishable, the probability function in the observation plane is
$P^{(1)}(x,t)$ $=$ $\langle|A_{S1x}+A_{S2x}|^{2}\rangle$ (103)
$=$ $\langle|e^{i\varphi_{j}}K_{S1x}+e^{i\varphi_{j}}K_{S2x}|^{2}\rangle$
@@
where $\varphi_{j}$ is the initial phase of the $j$th detected particle, $K_{S1x}$ is the particle’s Feynman propagator from S to $x$ in the observation plane via P1, the meanings of other symbols are defined similarly.
Particle’s Feynman propagator can be calculated in the same method as the one for photon and has simpler form for free particles Feynman et al. (2010)
@@
where $m$ is the mass of particle, $\hbar=h/(2\pi)$, $(x_{a},t_{a})$ and $(x_{b},t_{b})$ are the starting and ending space-time coordinates, respectively. For simplicity, one-dimension is considered.
Assuming S is in the symmetrical position of P1 and P2, there is no path difference from S to two pinholes. The path difference is introduced by the pinholes to the observation plane. Substituting Eq. (104) into Eq. (103), it is straightforward to have one-dimension spatial first-order interference pattern in the Fraunhofer range,
@@
in which paraxial approximation has been employed, $\lambda_{D}=h/(mv)$ is the de Broglie wavelength of particle, $v$ is the speed of atom and the speed of all atoms are assumed to be the same, $d$ is the distance between two pinholes, and $L$ is the distance between the pinhole and detection planes.
There is first-order interference pattern when these two different paths are indistinguishable for particles, no matter these particles are fermions or bosons, in cold atom state or BEC state. Just like there is no difference for the first-order interference pattern of photons in a Young’s double-slit interferometer for thermal and laser light. The initial phases of particles does not contribute to the first-order interference pattern of one beam.

### v2

@@ V Third- and higher-order interference of light
The third- and higher-order interference of light is a natural generalization of the second-order interference of light. The method introduced above can also be employed to calculate the third- and higher-order interference of light.
The history of the third- and higher-order interference dates back to 1963 when Glauber gave a unified definition of $N$th-order coherence function of light Glauber (1963a, c). Greenberg et al. generalized entangled photon pair to three-photon entanglement Greenberger et al. (1989, 1990). It was difficult to generate entangled three-photon directly then. Bouwmeester et al. employed two entangled photon pairs to generate entangled three-photon by post-selection Bouwmeester et al. (1999). Mitchell et al. employed a laser photon and an entangled photon pair to generate entangled three-photon and observed three-photon subwavelength interference Mitchell et al. (2004). In 2010, Hübel et al. employed cascade periodically poled nonlinear crystals to generate entangled three-photon directly Hübel et al. (2010). Based on the work of Bouwmeester et al. Bouwmeester et al. (1999), Walther et al. observed four-photon subwavelength interference with two entangled photon pairs Walther et al. (2004) and Pan’s group observed as high as 12-photon subwavelength interference Zhao et al. (2004); Pan et al. (2012); Zhong et al. (2018).
Ghost imaging was originally realized with entangled photon pairs Pittman et al. (1995) and later with thermal light Gatti et al. (2004); Cheng and Han (2004); Cai and Zhu (2005); Valencia et al. (2005) via the second-order interference of light. Ghost imaging with the third- and higher-order interference of light was discussed for entangled three-photon Wen et al. (2007) and thermal light Ou and Kuang (2007); Bai and Han (2007); Agafonov et al. (2008); Zhou et al. (2010); Cao et al. (2008); Chan et al. (2009); Liu et al. (2009).
In the following, third-order interference of light is taken as an example to show how to employ Feyman’s path integral to calculate the third- and high-order interference of light.
@@ V.1 Third-order interference of one light beam
Figure 27 shows the scheme for the third-order HBT interferometer by analogy of the HBT interferometer shown in Fig. 11. The light emitted by a light source, S, is split into two via a 1:2 non-polarizing beam splitter, BS1. The transmitted light beam is split into two via a 1:1 non-polarizing beam splitter, BS2. The intensities detected by D1, D2, and D3 are equal. TCC is short for three-photon coincidence count detection system and usually written as CC for convenience. The distance between the source and detector planes are equal.
Figure 27: Third-order HBT interferometer. TCC is short for three-photon coincidence count and is usually written as CC for convenience. BS1 is a 1:2 non-polarizing beam splitter. The meanings of other symbols are the same as the ones in a usual HBT interferometer.
Assume S is a thermal light source in Fig. 27. There are six different paths to trigger a three-photon coincidence count in D1, D2, and D3. Figure 28 shows the schematic diagram of the six different paths. Fig. 28(a) shows the path for that photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3. The corresponding three photon probability amplitude is $A_{a1,b2,c3}$. Other five paths are defined similarly. If all the six different paths are indistinguishable, the three-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{6}}(A_{a1,b2,c3}+A_{a1,b3,c2}+A_{a2,b1,c3}$
$+A_{a2,b3,c1}+A_{a3,b1,c2}+A_{a3,b2,c1})|^{2}\rangle.$
@@
Figure 28: Six different paths to trigger a three-photon coincidence count. (a) shows the path for photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3. Other five paths are defined similarly.
With the same method as the one for thermal light in a HBT interferometer in Sect. IV.1.1, the spatial three-photon probability function is Liu and Shih (2009)
@@
where paraxial approximation and one-dimension case were assumed, $\lambda$ is the wavelength of thermal light and single-frequency light is employed, $d$ is the length of the one-dimension thermal light source, $L$ is the distance between the source and detector planes. Figure 29(a) shows the simulated spatial third-order coherence function of thermal light in a third-order HBT interferometer. Figure 29(b) shows the measured three-photon coincidence counts vs. the transverse distances between these three detectors. The theoretical and experimental results are consistent with each other.
Figure 29: Simulated spatial third-order coherence function Liu and Shih (2009) and measured three-photon coincidence counts Zhou et al. (2010) of pseudothermal light in a third-order HBT interferometer. $x1-x3$ in (a) and $x_{13}$ in (b) represent transverse position difference of D1 and D3. Similar definition holds for $x1-x2$ in (a) and $x_{23}$ in (b). Triple C.C. in (b) is three-photon coincidence counts, which is proportional to the normalized third-order coherence function in (a).
Temporal three-photon probability function can be calculated in the same way Zhou et al. (2010),
@@
The degree of third order coherence, $g^{(3)}(0)$, of thermal light equals 6, which means that the probability of detecting three photons when these three detectors are in the same space-time coordinate is 6 times of the one when these three detectors are in different coherence volumes. It is defined as three-photon bunching of thermal light. The degree of $N$th-oder coherence of thermal light equals $N!$ Liu and Shih (2009). Three-photon superbunching can be observed with superbunching pseudothermal light Zhou et al. (2019a).
The third-order interference of single-mode continuous-wave laser light in a third-order HBT interferometer is simple. All the six different paths shown in Fig. 28 to trigger a three-photon coincidence count are identical. The superposition of these six probability amplitudes will not give three-photon interference. The reason is the same as the one for there is no two-photon bunching for single-mode continuous-wave laser light in a HBT interferometer. No three-photon bunching for single-mode continuous-wave laser light exists in the third-order HBT interferometer.
@@ V.2 Third-order interference of two independent light beams
Figure 30 shows the scheme for the third-order interference of two independent light beams. Assume S1 is a single-photon source and S2 is a single-mode continuous-wave laser. There are two different ways to trigger a three-photon coincidence count in Fig. 30. The first way is that one photon is emitted by S1 and two photons are emitted by S2. The probability is $3I_{1}I_{2}^{2}/(3I_{1}I_{2}^{2}+I_{2}^{3})$, where $I_{1}$ and $I_{2}$ are the intensities of light beams emitted by S1 and S2, respectively. There are three different paths and the corresponding three-photon probability amplitudes are $A_{11,22,23}$, $A_{12,21,23}$, and $A_{13,21,22}$, respectively. The second way is these three photons are all emitted by S2. The probability is $I_{2}^{3}/(I_{1}I_{2}^{2}+I_{2}^{3})$. There is one path and the corresponding three-photon probability amplitude is $A_{21,22,23}$.
Figure 30: Scheme for the third-order interference of two independent light beams. S1 and S2 are two light sources. D1, D2 and D3 are three single-photon detectors. The output signals of D1, D2 and D3 are sent into a three-photon coincidence count detection system, which is not shown.
@@
The initial phases of photons emitted by the single-photon source, S1, are random. The initial phases of photons emitted by the single-mode continuous-wave laser, S2, are identical within one coherence volume. Equation (96) can be simplified as
@@
There is third-order interference pattern. The visibility of the third-order interference pattern is dependent on the ratio between the intensities of light emitted by the single-photon source and single-mode continuous-wave laser.
@@ V.3 Third-order interference of multiple independent light beams
Figure 31 shows the scheme for the third-order interference of three independent light beams. Assume S1, S2, and S3 are three identical and independent single-photon sources. The intensities, polarizations, and frequency spectra of the photons emitted by these three sources are the same. There are six different paths to trigger a three-photon coincidence count. The corresponding three-photon probability amplitudes are $A_{11,22,33}$, $A_{11,23,31}$, $A_{12,21,33}$, $A_{12,23,31}$, $A_{13,22,31}$, and $A_{13,21,32}$, where $A_{11,22,33}$ is the probability amplitude for that the photon emitted by S1 is detected by D1, the photon emitted by S2 is detected by D2, and the photon emitted by S3 is detected by D3. The meanings of other symbols are defined similarly.
Figure 31: Scheme for the third-order interference of three independent light beams. The meanings of the symbols are similar as the ones in Fig. 30.
If all the six different paths are indistinguishable, three-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{6}}(A_{11,22,33}+A_{11,23,31}+A_{12,21,33}$
$+A_{12,23,31}+A_{13,22,31}+A_{13,21,32})|^{2}\rangle$
$\propto$ $\langle|e^{i(\varphi_{1}+\varphi_{2}+\varphi_{3})}(K_{11}K_{22}K_{33}+K_{11}K_{23}K_{31}$
$+K_{12}K_{21}K_{33}+K_{12}K_{23}K_{31}$
$+K_{13}K_{22}K_{31}+K_{13}K_{21}K_{32})|^{2}\rangle,$
@@
With the same method and approximations above, one-dimension spatial three-photon probability function can be simplified as
@@
Figure 32 shows the third-order interference pattern given by Eq. (V.3). The periodic structures can be observed in $x_{1}$, $x_{2}$, and $x_{1}-x_{2}$ directions.
Figure 32: Third-order interference pattern of three independent single-photon states. $x_{1}$ and $x_{2}$ are the transverse positions of D1 and D2, respectively.
Similar method can be employed to calculate the third-order interference other types of light beams and higher-order interference of light. In fact, all optical coherence problems can be calculated in the same method introduced above.
@@ VI Quantum atomic coherence theory based on Feynman’s path integral
It is straightforward to generalize quantum optical coherence theory based on Feynman’s path integral to discuss the coherence of massive particles. The history of atomic coherence traces back to de Broglie in 1924 when he proposed matter wave assumption De Broglie (1924) by analogy of Planck’s Planck (1901) and Einstein’s Einstein (1905) quanta assumption of light. The de Broglie wavelength of matter wave equals
@@
De Broglie’s assumption was soon verified by Davisson et al. and Thomson et al. with diffraction pattern of electrons Davisson and Germer (1927); Thomson and Reid (1927). Due to the very short de Broglie wavelength of matter wave, it was difficult to observe the first-order interference pattern like the one of light then. In 1953, Marton et al. observed first-order interference pattern of electrons in a Mach-Zehnder interferometer Marton et al. (1953). Then the first-order interference pattern of neutron Maier-Leibnitz and Springer (1962), atom Keith et al. (1991), molecule Arndt et al. (1999) were also reported. The first-order interference pattern of two independent Bose Einstein Condensations (BEC) was observed Andrews et al. (1997).
The first experiment of the second-order interference with matter wave was done by Yasuda and Shimizu, in which they observed two-atom bunching with cold atoms Yasuda and Shimizu (1996). Burt et al. found that the ratio between the degrees of third-order coherence of cold atoms and BEC equals $7.4\pm 2.6$ Burt et al. (1997). The second-order interference of fermions was reported by several groups Henny et al. (1999); Oliver et al. (1999); Iannuzzi et al. (2006); Rom et al. (2006); Bocquillon et al. (2013).
Quantum atomic coherence theory is mainly based on Glauber’s quantum optical coherence theory Adams et al. (1994); Meystre (2001); Cronin et al. (2009). Feynman himself introduced how to employ Fenman’s path integral to calculate the first-order interference of electrons in a Young’s double-slit interferometer Feynman et al. (2010, 2011). Storey and Cohen-Tannoudji provided an introduction of how to employ Feynman’s path integral to calculate the first-order interference of atoms Storey and Cohen-Tannoudji (1994). Employing Feynman’s path integral to calculate the first-order interference of matter wave can be found in Refs. Deng (2006, 2008); Gondran and Gondran (2005), too.
This section will focus on how to employ Feynman’s path integral to calculate the first-, second, and higher-order interference of matter waves, with special attentions on the analogy between cold atoms (BEC) and thermal light (laser light).
@@ VI.1.1 First-order interference of one matter wave beam
Figure 33 shows the scheme for the first-order interference of one matter wave beam. S is a particle source. P1 and P2 are two pinholes with equal size. O is the observation plane. For simplicity, free atoms are assumed here and in the following discussions.
Figure 33: Scheme for the first-order interference of one matter wave beam. S is a particle source. P1 and P2 are two pinholes. O is the observation plane.
There are two different paths to trigger a particle detection event in the observation plane. The first path is the particle going through P1. The probability is 1/2 and the probability amplitude is $A_{S1x}$. The second path is the particle going through P2. The probability is 1/2 and the probability amplitude is $A_{S2x}$. If these two different paths are indistinguishable, the probability function of particles in the observation plane is
$P^{(1)}(x,t)$ $=$ $\langle|\frac{1}{\sqrt{2}}(A_{S1x}+A_{S2x})|^{2}\rangle$ (103)
$\propto$ $\langle|e^{i\varphi_{j}}K_{S1x}+e^{i\varphi_{j}}K_{S2x}|^{2}\rangle$
@@
where $\varphi_{j}$ is the initial phase of the $j$th detected particle, $K_{S1x}$ is the particle’s Feynman propagator from S to $x$ via P1, the meanings of other symbols are defined similarly.
Massive particle’s Feynman propagator can be calculated in the same method as the one for photon and has a simple form for free particles Feynman et al. (2010)
@@
where $m$ is the mass of particle, $\hbar=h/(2\pi)$, $(x_{a},t_{a})$ and $(x_{b},t_{b})$ are the starting and ending space-time coordinates of the particle, respectively. For simplicity, one-dimension case is assumed.
Assuming S is at the symmetrical position of P1 and P2, there is no path difference from S to two pinholes. The path difference is introduced by the distances from the pinholes to the observation plane. Substituting Eq. (104) into Eq. (103), it is straightforward to have one-dimension spatial first-order interference pattern in the Fraunhofer range,
@@
in which paraxial approximation has been employed, $\lambda_{D}=h/(mv)$ is the de Broglie wavelength of the particle, $v$ is the speed of the particle and the speed of all the particles are assumed to be the same, $d$ is the distance between these two pinholes, and $L$ is the distance between the pinhole and observation planes.
There is first-order interference pattern when these two different paths are indistinguishable. The patterns are the same for fermions and bosons. Just like there is no difference for the first-order interference patterns of thermal and laser light in a Young’s double-slit interferometer, the first-order interference patterns of bosons in thermal state or BEC state in a Young’s double-slit interferometer are the same. The initial phases of the particles do not influence the first-order interference pattern of one beam.